OpenStudy (aaronandyson):

Prove That sec.cosec = tan + cot + 2 @welshfella @Loser66

OpenStudy (aaronandyson):

@iGreen

OpenStudy (aaronandyson):

@misty1212

OpenStudy (loser66):

play from the right handside. Breaking tan = sin/cos, cot = cos/sin. Tip: 2 = \(\dfrac{2sincos}{sincos}\)

OpenStudy (aaronandyson):

What?

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

\[\sec(x)\csc(x) = \tan(x) + \cot(x) + 2\] like that?

OpenStudy (aaronandyson):

yes

OpenStudy (aaronandyson):

@misty1212 @Loser66

OpenStudy (aaronandyson):

@misty1212 @Loser66 @mathstudent55 @mathmate @mathmath333

OpenStudy (aaronandyson):

@triciaal

OpenStudy (aaronandyson):

@welshfella

OpenStudy (aaronandyson):

@misty1212 @Loser66 @mathstudent55 @mathmate @mathmath333

OpenStudy (loser66):

I gave you all. you just work on it. Show me your attempt, please.

OpenStudy (aaronandyson):

I can't I'm confused what to do

OpenStudy (loser66):

If I do it for you, you NEVER know how to do. Try once. Follow what I guided.

OpenStudy (aaronandyson):

sin/cos + cos/sin + 2 now?

OpenStudy (aaronandyson):

@Loser66

OpenStudy (loser66):

where is my tip for 2?

OpenStudy (aaronandyson):

I can't understand your tip2..

OpenStudy (loser66):

that is replace 2 by 2sincos/sincos

OpenStudy (aaronandyson):

sin/cos + cos/sin + 2sincos/sincos??

OpenStudy (loser66):